OpenStudy (anonymous):

please help I don't know how to do this Factor this expression completely. x2 - r2 I will medal

OpenStudy (anonymous):

(x+r)(x-r)

OpenStudy (anonymous):

can you explain this to me?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay then..

OpenStudy (anonymous):

2

OpenStudy (mrnood):

to factor something you need to break it up into components that multiply togjether to make the original. so you can see that x^2 must break into x * x

OpenStudy (anonymous):

oh okay that makes sense thank you

OpenStudy (mrnood):

so the factors must be (x+a)(x+b) if you multiply that out you will get x^2 +ax +bx + b^2 if you compare that with the original you can see that there are NO terms with x in them therefore a=-b (so that they cancel out) if you compare with te original again you can see that the b^2 term equates to r^2 therefore b = r so we know b= r and a=-b so a= -r so the factors are (x+r)(x-r)

OpenStudy (anonymous):

@MrNood Factor this expression completely. 2x^2 + 11^x + 15 (2*x+x)+(11*x+15)?

OpenStudy (mrnood):

you can check that by expanding again: (x+r)(x-r) = x^2 +rx -rx -r^2

OpenStudy (mrnood):

so we are looking for factors tha tproduce a term with 2x^2 this can ONLY be 2x * x so our factors are (2x+a)(x+b) try to expand that pair and see what you get.... I'll help if oyu do that...

OpenStudy (anonymous):

expand (2x+a)(x+b)?

OpenStudy (mrnood):

yes - multiply them out completely

OpenStudy (anonymous):

2xa*xb?

OpenStudy (anonymous):

or 2x^2ab

OpenStudy (mrnood):

no use 'FOIL' method? you must multiply both terms in each bracket by both in the other

OpenStudy (anonymous):

idk what FOIL method is or I can't recall it

OpenStudy (anonymous):

2x^2ab*2x^2ab? is it that?

OpenStudy (anonymous):

??? @MrNood

OpenStudy (mrnood):

 FOIL is 'First, Outer, Inner, Last' Multiply first pair in each bracket then outer pair then inner pair then last pair